# Extreme Nuclear Transients Resulting from the Tidal Disruption of Intermediate Mass Stars

### v1

@@ Acknowledgement
https://www.cosmos.esa.int/gaia. The Gaia archive website is https://archives.esac.esa.int/gaia. This work is also based in part on observations obtained with the Samuel Oschin Telescope 48-inch and the 60-inch Telescope at the Palomar Observatory as part of the Zwicky Transient Facility project. ZTF is supported by the National Science Foundation under Grant No. AST-2034437 and a collaboration including Caltech, IPAC, the Weizmann Institute for Science, the Oskar Klein Center at Stockholm University, the University of Maryland, Deutsches Elektronen-Synchrotron and Humboldt University, the TANGO Consortium of Taiwan, the University of Wisconsin at Milwaukee, Trinity College Dublin, Lawrence Livermore National Laboratories, and IN2P3, France. Operations are conducted by COO, IPAC, and UW.
@@
Figures S1 through S9
Table S1 through S3
References 55 through 198
@@
Sample Selection.
We select our initial sample of extreme nuclear transients (ENTs) from the Gaia Alerts (?) transient stream. Our criteria for inclusion were designed to select smooth, luminous, and long-lived events. We had five selection criteria: (1) a flare of mag over the pre-flare baseline, (2) an observed timescale of that flare of year, (3) a flare that was smooth as defined by having a monotonic flare profile and no strong short-term variability (10% excess variability) during the flare, (4) a flare with a peak luminosity of erg s-1, and (5) a source without radio and/or gamma-ray detections that would suggest a jetted AGN. This selection resulted in two sources: Gaia16aaw and Gaia18cdj.
@@
Thus, the three transients we study in detail are Gaia16aaw, Gaia18cdj, and ZTF20abrbeie. Gaia16aaw, (04:11:57.000, 42:05:30.80), was discovered on 2016 January 23.5 by the Gaia Alerts team (?). The discovery was announced publicly on the Transient Name Server (TNS) and given the identification AT2016dbs111https://www.wis-tns.org/object/2016dbs (?). Gaia18cdj, (02:09:48.140, 42:04:37.02), was discovered on 2018 August 12.1 by the Gaia Alerts team (?) and announced on TNS with the identification AT2018fbb222https://www.wis-tns.org/object/2018fbb (?). ZTF20abrbeie, (21:13:48.408, 27:25:50.48), was discovered on 2021 April 13.5 by ZTF. It was announced to TNS with the identification AT2021lwx333https://www.wis-tns.org/object/2021lwx (?). Color images created using aplpy (?) from images taken by either the Dark Energy Survey (DES; for Gaia16aaw and Gaia18cdj) (?) or the Panoramic Survey Telescope and Rapid Response System (Pan-STARRS; for AT2021lwx) (?) are shown in Figure S1.
As ZTF20abrbeie has had previous papers published on its evolution under the name AT2021lwx (?, ?), we will continue to use that name in this manuscript to avoid confusion. For the two Gaia sources, we elect to use their survey names in this manuscript.
Archival and Transient Photometry.
We first searched for available archival photometry to constrain the evolution of these objects prior to the ENT flare. For Gaia16aaw and Gaia18cdj, we obtained -band photometry from the CRTS. As the typical signal-to-noise ratio (S/N) per CRTS epoch was low, we binned these data in monthly bins to search for pre-flare variability. These data are shown in Figure 1 as tan squares. The sources are weakly detected and show no significant evidence for strong variability prior to the flares.
@@
For AT2021lwx, there is no persistent source detected in the NEOWISE single exposure catalog, consistent with the apparent lack of a host galaxy in Pan-STARRS imaging (?, ?). We therefore performed aperture photometry on the NEOWISE images to obtain our WISE light curves. To avoid contamination by nearby, bright stars, we used a 4” radius aperture with an aperture correction estimated from the computed WISE curve of growth444https://wise2.ipac.caltech.edu/docs/release/allsky/expsup/sec4_4c.html. The local background was estimated using a sigma-clipped median within an annulus around the target. Despite the small source aperture, the band retained a low level of residual flux from the nearby, bright star. To mitigate this, we subtracted the mean pre-flare flux from the light curves and added the pre-flare scatter in quadrature with the estimated flux uncertainties for each epoch. The WISE light curves are shown in Figure 1 as olive and brown diamonds for and respectively. Given the lack of a pre-flare detection, we only show NEOWISE data for AT2021lwx.
@@
Finally, we acquired light curves from the discovering survey for each source. For Gaia16aaw and Gaia18cdj, this was -band photometry from Gaia, obtained through the Gaia Alerts service. We estimated the uncertainties on this photometry following (?) and binned visits within 5 days for subsequent analyses. The Gaia light curves are shown in Figure 1 as green hexagons. For AT2021lwx, we procured ZTF photometry in the and bands from the Alerce broker (?). During the rise and through peak, we stacked the data 5-day bins, switching to monthly bins after the last seasonal break to better sample the decline. These data are shown in Figure 1 as teal and red octagons.
Offset from Host-Galaxy Center.
To constrain the offset of the ENTs from the center of their host galaxies, we aligned the images and then measured the offset between the positions before and during the transient. We elected to measure the relative position before and during transient emission rather than absolute positions to avoid uncertainty in the distortion terms that going to a full world coordinate solution would induce. We first used a modified version of the ISIS image subtraction package (?, ?) interpolation function which matches sources identified by sextractor (?, ?) to align the DES images at the location of Gaia16aaw and Gaia18cdj. We can not measure the offset for AT2021lwx as the host galaxy was not detected in any archival images. For the Gaia sources, we first retrieved the DES images from the NOIRlab image servers. Because the image cutouts are not aligned, we first did a rough alignment of all retrieved cutouts including all DES filters. We then removed image cutouts with too small an area to have a sufficient number of alignment stars, images where the source was near the edge, and images where the interpolation did not converge. We then trimmed the images to the intersecting area and re-interpolated the images. This left 20 images for Gaia16aaw and 8 images for Gaia18cdj.
Next, we used photutils (?) to determine the centroid of the host galaxy before the transient and then again when each transient was near peak. For Gaia16aaw, the DES images that pass our cuts are from Dec 2013, Jan 2014, Jan 2016, Feb 2016, Nov 2016, Dec 2016, Dec 2017, and Jan 2018. The available Gaia light curve begins in Oct 2014 and it appears that the host+transient emission may already be brighter than the host alone because it was mag brighter in 2014 than in 2023, indicating that the source was already on the rise. Thus we use the 8 DES images that passed our cuts from Dec 2013 and Jan 2014 for our host galaxy image without the transient flux. The light curve of Gaia16aaw peaks on 2016 March 28 and fades by only mag by Dec 2016. We use the 9 DES images from 2016 to centroid the transient as these are dominated by the transient emission. We average the centroids from the host galaxy and transient epochs and use the standard deviation of these measurements as an estimate of the statistical error. Finally, we use 10 stars of comparable brightness to Gaia16aaw and centroid those stars in all our interpolated DES images. We then take the median of the standard deviation of each star’s positions as an estimate of our systematic uncertainty. This yielded an offset from the host galaxy’s center of which corresponds to a physical offset of kpc at the distance of Gaia16aaw.
We then repeat this process for Gaia18cdj. Gaia18cdj has DES epochs from Oct 2013, Jan 2014, Oct 2016, Nov 2016, Nov 2018, and Dec 2018. The light curve of Gaia18cdj peaks on 2018 October 24. There is no detectable additional transient flux from Nov 2016 and earlier so we use the 5 DES images that pass our cuts during these times for the host galaxy and the 3 DES images in 2018 to centroid the transient position. We then follow a similar procedure to Gaia16aaw including 10 comparison stars. This yielded an offset from the host galaxy’s center of which corresponds to a physical distance offset of kpc at the distance of Gaia18cdj.
Follow-up Spectroscopy.
After identification of the two Gaia sources as ENTs, we obtained follow-up spectra. For Gaia16aaw, we obtained an optical spectrum on MJD 59210.1 (854 rest-frame days after peak) using the Inamori-Magellan Areal Camera and Spectrograph (IMACS; (?)) on the 6.5-m Magellan Baade telescope. For Gaia18cdj, we obtained an optical spectrum on MJD 59168.2 (388 rest-frame days after peak) using the Goodman High Throughput Spectrograph (?) on the Southern Astrophysical Research telescope. These spectra were reduced and calibrated with standard IRAF (?, ?) procedures such as bias subtraction, flat-fielding, 1-D spectroscopic extraction, and wavelength calibration. For AT2021lwx, we obtained data for the first observer-frame optical spectra presented in (?) from the Keck Observatory Archive and reduced it using PypeIt (?). The flux calibrations were initially performed using standard star spectra and then scaled to match concurrent Gaia photometry for Gaia16aaw and Gaia18cdj and ZTF -band photometry for AT2021lwx. These spectra are shown in Figure S2. We additionally obtained the follow-up optical spectra from (?) and (?) for AT2021lwx, and again used ZTF -band photometry to refine the flux calibration.
From these spectra, we can estimate the redshifts of the sources. For AT2021lwx, we confirm the redshift of (?, ?). For Gaia18cdj there is a clear Mg II absorption doublet that places the source at . The fact that there is Mg II emission at the same redshift confirms that this absorption feature is not a foreground absorber. The spectrum of Gaia16aaw does not exhibit any strong narrow emission or absorption features, and the redder portions of the spectrum are noisy due to night sky lines and the faint magnitude of the source at the time the spectrum was taken ( AB mag). However, there is a strong broad feature that we interpret as Mg II, placing the source at . Our identification of this feature as Mg II is supported by the broad Fe feature just blueward of the Mg II feature. Such a feature is seen in the spectra of both Gaia18cdj and Gaia16aaw.
@@
Stellar Mass and Star Formation Rate.
The host galaxies of Gaia16aaw and Gaia18cdj are readily detected in archival surveys, so we model their properties. We used the Code Investigating GALaxy Emission (CIGALE; (?, ?, ?)) because it allows for simultaneous contributions from AGN activity and stellar emission. We fit the optical () data from co-adding the pre-flare epochs of DES imaging, near-infrared () data from the VISTA Hemisphere Survey (?), and mid-infrared () data from the AllWISE catalog (?).
We used a delayed star formation history, the (?) stellar population models, a Salpeter initial mass function (?), a (?) extinction law with RV = 3.1, and the SKIRTOR AGN model (?, ?). From our CIGALE fits, we find that the host galaxy of Gaia16aaw has a stellar mass of M M⊙, an age of Gyr, and a star formation rate of SFR M⊙ yr-1. For the host of Gaia18cdj the values are M M⊙, an age of Gyr, and a SFR M⊙ yr-1. We used synthetic photometry computed from the best-fitting host-galaxy models to subtract the host-galaxy fluxes from photometry as needed to isolate the transient flux.
@@
We further examine the SFRs of these ENT host galaxies using narrow emission lines and standard scaling relations. Gaia16aaw shows an emission feature near [O II], although there is significant contamination from night sky lines. Assuming that this line is indeed [O II], we derive a flux of erg s-1 cm-2. At the distance of Gaia16aaw and correcting for the mag extinction from the CIGALE fits, this corresponds to a luminosity of erg s-1. From typical scaling relations (?, ?), such a luminosity implies a SFR of M⊙ yr-1, roughly 2 below the estimate from the CIGALE fits.
Gaia18cdj does not show an [O II] emission line, so we calculated an upper-limit on the [O II] luminosity following the procedure of (?). We assumed a line width of 300 km s-1 and obtained 3 flux limits as
@@
where is the continuum flux at wavelength , is the RMS scatter about the normalized continuum, is the width of the line profile, and is the pixel scale of the spectrum. Correcting for the extinction from the CIGALE fits, this yields a luminosity upper-limit of erg s-1, implying an SFR of M⊙ yr-1 (?, ?). While the SFRs estimated from spectroscopic scaling relations are low relative to the estimates from CIGALE, we note (1) that these lines are in regions of the spectra with lower S/N and (2) that the SFRs estimated from the rest-frame UV absolute magnitudes of mag (?) are consistent with the CIGALE estimates.
The host galaxy of AT2021lwx is undetected even in deep pre-outburst optical and IR imaging, precluding any detailed analysis of its properties. However, previous works have estimated upper limits on the stellar mass and star formation rate (SFR). From the Pan-STARRS upper limits and assuming a mass-to-light ratio of 2, (?) compute a mass limit of M⊙. Since there is considerable scatter in the mass-to-light ratio for the bluer rest-frame bands being probed by the Pan-STARRS data, up to a (?), we will instead adopt a slightly more permissive limit of M M⊙ here. From limits on the [O II] luminosity in their follow-up spectra, (?) calculate an upper limit on the SFR of M⊙ yr-1, which we confirm using deep follow-up spectra.
The stellar masses and SFRs from CIGALE for the hosts of Gaia16aaw and Gaia18cdj and the estimated limits for AT2021lwx are shown in Figure S3. Along with these three ENTs, we show host properties for a sample of tidal disruption events (?) in blue and a sample of ambiguous nuclear transients (?, ?) in gold. We additionally show two broader galaxy samples in gray, one local sample from SDSS (?) and the MPA-JHU catalog (?) and a sample of galaxies at from Cosmic Evolution Survey 2020 data release (COSMOS; (?)). We have also indicated the SFRs expected for the star-forming main sequence at (?) with the dashed line. In addition to the individual source host properties, we show kernel density estimates (KDE) computed using scipy.stats.gaussian_kde and Scott’s Rule to model the underlying distributions.
The stellar masses and SFRs of the ENT hosts are generally higher than any of the hosts in comparison samples. As compared to the TDE sample, the stellar masses are more than an order of magnitude larger than the typical TDE host and the SFRs are several orders of magnitude higher. When compared to the ANTs, the difference is less stark in mass, although the SFRs are still much higher. Finally, as compared to the local galaxy population, the host masses are consistent with the massive end of the blue sequence, but with SFRs nearly two orders of magnitude higher than the expected SFR for this mass. However, this is expected as these galaxies reside at . When compared to the location of the star-forming main sequence at this redshift as probed by the COSMOS data, the difference is less dramatic. In particular, when compared to the expected SFR (?) given the star-forming main sequence at the redshift and stellar mass of Gaia16aaw and Gaia18cdj, they lie 0.45 and 0.36 dex above the relation respectively. As compared to the typical dispersion about the star-forming main sequence of dex (?, ?, ?, ?), these hosts appear high, but consistent with, a typical massive star-forming galaxy at . The masses of the hosts for Gaia16aaw and Gaia18cdj are quite high for , far more massive than the typical galaxy at that redshift.
Using the scaling relationship between stellar mass and SMBH mass (?, ?), with a typical scatter of dex, we can estimate the central black hole masses. For the hosts of Gaia16aaw and Gaia18cdj, which have similar stellar masses, we find a M M⊙. Taking the limit on stellar mass for AT2021lwx, we find M M⊙, which is consistent with previous estimates (?, ?).
Presence of AGN Activity.
We can also assess the presence of AGN activity through the WISE color and associated selection criteria (?, ?, ?). Gaia16aaw has a color of Vega mag. This is redder than the mag threshold for local galaxies and the mag color expected for star-forming galaxies at this redshift. However, the position of AGNs within this color space changes with redshift. For the AGN template of (?) at z = 1.03, the color is 1.41 Vega mag, suggesting that while the host of Gaia16aaw likely hosts an AGN, it may not be dominating the MIR emission.
While the host of AT2021lwx is undetected and therefore we cannot use CIGALE or typical color selections to constrain pre-flare AGN activity, the archival photometry and follow-up spectra can be used to place constraints. AT2021lwx is undetected in Pan-STARRS with mag (?). If we assume that the corresponding rest-frame UV emission (Å) is solely tracing light from an AGN, this places a limit on the AGN luminosity of erg s-1, ruling out a quasar. Additionally, the spectra in (?) and (?) show no signs of strong narrow emission lines from AGN activity. Using Eq. (1), we computed an upper limit on the [O III] luminosity assuming a line width of 1000 km s-1. This yields a 3 upper-limit of erg s-1 which also rules out a luminous quasar but is consistent with a Seyfert (e.g., (?, ?, ?)).
Finally, we can use archival X-ray data to probe pre-flare AGN activity in these galaxies. From ROSAT All-Sky Survey (RASS) data we calculate count rates for the host galaxies of our ENTs and convert them to keV fluxes assuming a photon index of and the Galactic foreground column density (?). Gaia16aaw is detected in ROSAT, with a keV luminosity of erg s-1, significantly higher than the erg s-1 expected from the high SFR (?, ?), thus confirming the presence of an AGN. Gaia18cdj is undetected, with a 3 upper-limit of erg s-1. Gaia18cdj is also undetected in the later Swift XRT coverage, providing a deeper limit of erg s-1. This further rules out a strong quasar, but this limit is consistent with a weaker AGN. For AT2021lwx, the limit from RASS (using ; see the X-ray emission section of the supplementary text) is erg s-1, which does not constrain the pre-flare AGN activity.
Context Within the Transient Landscape
The corresponding peak absolute magnitudes, accounting for foreground extinction and applying a flat K-correction of , are M, M, and M for Gaia16aaw, Gaia18cdj, and AT2021lwx respectively. The uncertainties are only statistical errors from the Monte Carlo procedure. We also show the source ASASSN-15lh as a comparison object given its status as a similarly overluminous transient (?, ?). From the same procedure, applied to the ASAS-SN -band light curve, we estimate a peak absolute magnitude of M.
Figure 4 compares the ENTs to a number of different types of transients. In blue we show various classes of supernovae, ranging from the fast and faint Calcium-Rich transients (e.g., (?)) to the superluminous supernovae (SLSNe; (?)). The green boxes show transients related to stellar mergers and/or mass transfer (e.g., (?, ?)). In red shades, we show transients powered by accretion onto SMBHs, including tidal disruption events (?). We have also estimated the absolute magnitude and characteristic timescale range for the growing class of ambiguous nuclear transients (?, ?, ?, ?, ?). While their connection to SMBH accretion is likely, we have elected to present them in a light shade of red to indicate the uncertainty regarding their physical origin. It is clear that the ENTs studied here lie at higher peak luminosities than any other source class and are among the longest-lived. The only other transients that rival the long characteristic timescales are the core-collapse supernovae, which are mag fainter, and some ANTs, which are still several magnitudes fainter at peak.
The ranges shown for the various classes of transients are broadly based on (?). Given the large increase in discoveries since that work, we have augmented the regions for certain classes, such as TDEs and SLSNe, to reflect the full range of properties better. Nevertheless, it is clear that these ENTs reside in a region currently unoccupied by other transient classes.
There is also a class of AGN flares discovered by CRTS (?) that bears some resemblance to the ENTs. Many of the CRTS AGN flares are luminous and long-lived, but they typically are not as smooth, with many showing short-term variability or long-timescale re-brightening episodes. Even for the apparently smooth flares, the low S/N per CRTS epoch can easily mask variability, unlike the comparatively higher S/N Gaia and ZTF data available for the ENTs in our sample. Since we cannot confirm their smooth nature, which is one of our ENT selection criteria, we do not consider them in the remainder of the manuscript.
Properties of the Flares
Spectral Energy Distribution.
We can go beyond just the colors of the sources and estimate their properties through blackbody fits to their time-evolving SEDs. Using Markov Chain Monte Carlo (MCMC) methods and a forward-modeling approach, we fit the available epochs of host-subtracted and foreground extinction-corrected multi-band photometry as a blackbody to obtain the bolometric luminosity, temperature, and effective radius for the ENTs. For Gaia16aaw, we fit three epochs of DES data, for Gaia18cdj we fit two epochs of DES data and one epoch of Swift photometry, and for AT2021lwx we fit six epochs of Swift data. For the DES data, we added a 2% uncertainty in quadrature with the photometric uncertainties before fitting. These fits are given in Table S1. We find that a blackbody model adequately describes the observed emission, with a median reduced of 2.3, 0.8, and 0.7 for Gaia16aaw, Gaia18cdj, and AT2021lwx respectively.
In Figure S5 we show the results for the ENT luminosity, radius, and temperature compared to a sample of well-observed TDEs (?), SLSNe-I from (?), and ANTs (?, ?). In terms of luminosity, the ENTs are clearly more luminous and longer-lived, as expected given their position in Fig. 4. Their decline slopes are flatter than all of the TDEs and SLSNe, but similar to some of the ANTs. Two transients, ASASSN-15lh (either a SLSN or TDE; (?, ?, ?)) and AT2019brs (an ANT; (?)) have similar peak luminosities, although the initial decline rate for ASASSN-15lh is significantly steeper and thus the integrated energy is lower.
The effective radii of our ENTs are significantly larger than the TDEs, but consistent with the SLSNe and ANT comparison samples. In terms of the radius evolution, the ENTs appear to show a slow decline in radius post-peak. This is consistent with the TDE and ANT evolution, but not with SLSNe whose effective radii increase with time. The effective temperatures of the ENTs lie between those of the SLSNe and TDE samples and are strikingly similar to the ANTs. The temperature of Gaia16aaw slowly cools after peak whereas Gaia18cdj slightly increases during the flare. Such behavior is more consistent with either TDEs or ANTs than the cool temperatures of SLSNe that drop dramatically after peak as the ejecta expand and cool (e.g., (?)).
For each of our sources, we also created a bolometric light curve by scaling the host-subtracted Gaia (for Gaia16aaw and Gaia18cdj) and ZTF (for AT2021lwx) light curves to match the bolometric luminosity estimated from the blackbody fits. For epochs between blackbody fits we linearly interpolated the bolometric correction. For data outside of this range we used a flat bolometric correction corresponding to either the first or last fitted blackbody epoch. This is similar to what we have done for previous events (?, ?, ?, ?) and gives a peak luminosity for AT2021lwx consistent with previous estimates (?, ?).
Figure 3 compares the ENT bolometric light curves to luminous nuclear transients of various classes. As in Fig. S5 the ENTs are generally the most luminous transients, although ASASSN-15lh is more luminous than Gaia16aaw at peak. It is clear that the ENTs decay more slowly than the majority of other transients.
Finally, we used the bolometric light curves to quantify the smoothness of the ENT flares. We used a Savitzky–Golay filter (?) with a window of 100 rest-frame days and a cubic polynomial to compute the long-term trend. We then normalized the bolometric light curves by this long-term trend and computed the RMS scatter as the characteristic fractional variability of the flare about the overall flare profile. This was 12% for Gaia16aaw, 9% for Gaia18cdj, and 6% for AT2021lwx. After subtracting the median fractional uncertainty of the bolometric light curve in quadrature (e.g., (?)), Gaia16aaw has an excess variability of 6% and both Gaia18cdj and AT2021lwx are consistent with noise. These confirm that the ENT flares are smooth compared to typical quasars (e.g., (?)).
Radiated Energy.
We also estimated the total radiated energy by fitting a Gaussian to the early-time rise and an exponential decline to the late-time decline. This allowed us to smoothly extrapolate to times without observational constraints using rise and decay slopes motivated by the existing data. We integrated the Gaussian fit before the first epoch of data and the exponential decline after the last epoch and added these energies to the energy computed by directly integrating the bolometric light curve. Given the high S/N data for these ENTs, the statistical uncertainties on the emitted energy from the bolometric light curve are small. Therefore, we conservatively estimate the uncertainty on the energies as half of the difference between the directly-integrated energy and the total energy from adding the fits to unobserved portions of the transient evolution. The total radiated energy for Gaia16aaw was erg, for Gaia18cdj it was erg, and for AT2021lwx the energy was erg. Unsurprisingly, the estimated total energies are not significantly higher than the directly-integrated energies since each of the ENTs has been observed late into their evolution. Assuming these flares are powered by accretion with an efficiency of 10%, the energies correspond to accreted masses of M⊙.
These energies are significantly higher than other known transients. Following the same procedure, we estimate a total emitted energy of erg for ASASSN-15lh, less than half that of Gaia16aaw, already the least energetic of the three flares studied here. The energetic transient PS1-10adi (?), with a total energy of erg also lies below these flares. These flares are roughly an order of magnitude more energetic than typical well-observed SLSNe (e.g., (?)), TDEs (e.g., (?)), and ANTs (e.g., (?)) and at least several times more energetic than the most energetic examples of each class.
Dust Covering Fraction.
We follow a similar procedure as (?). Briefly, we subtract the pre-flare emission from the and light curves and fit the host-subtracted and extinction-corrected WISE photometry as a blackbody. From the luminosity evolution in the IR and UV/optical, we fit for the peak luminosity in each band. The dust covering fraction (fc) is estimated as the ratio of the peak IR luminosity to the peak UV/optical luminosity (e.g., (?)).
Our blackbody fits to the NEOWISE data yield dust temperatures of K and effective radii of pc with the temperatures decreasing and the radii increasing in time, consistent with previous studies of hot dust in galactic nuclei (?, ?, ?). From the luminosity ratios, we find only a lower-limit of f for Gaia16aaw, as the peak of the IR luminosity is not observed. For Gaia18cdj we find f and for AT2021lwx we find f, where the uncertainties are estimated from Monte Carlo resampling.
These covering fractions are shown in Figure S6, where they are compared to previously estimated covering fractions for TDEs (?, ?) and ANTs (?). The ENT dust covering fractions are much higher than for optically-selected TDEs, but very similar to both ANTs and the minimum covering fraction estimates for the IR-selected TDEs (?). The ENT covering fractions are also remarkably consistent with the typical covering fraction of AGNs with (e.g., (?)). The nuclei of the ENT hosts are clearly dusty and likely harbor an AGN-like “dusty torus”, unsurprising given the existing evidence for AGN activity in the host galaxies of Gaia16aaw and Gaia18cdj.
@@
Anisotropies in the dust distribution can cause the ratio of IR to UV/optical luminosity to underestimate low covering factors and overestimate high covering factors. While the emitting geometry of these ENTs is unknown, we apply the corrections of (?) to examine their effects. Using Table 1 of (?) and assuming an aligned disk and torus with , their minimum computed optical depth, we find covering fractions of f, f, and f for Gaia16aaw, Gaia18cdj, and AT2021lwx respectively. While these corrections increase the covering fraction estimates for our ENTs, their position relative to the TDE and ANT samples as well as the typical AGN trend in Fig. S6 is not dramatically different.
@@
Finally, we used JAVELIN (?, ?) to compute temporal lags between the optical and IR flares, employing a top hat smoothing function. For AT2021lwx, as the source is undetected pre-flare, we added artificial optical measurements at the time of the NEOWISE observations with zero flux and uncertainties equal to the first optical flux uncertainty. This was necessary to avoid significantly negative fluxes in the model optical light curve at early times. These fits are shown in Figure S7. We find rest-frame lags of 265 days for Gaia16aaw, 150 days for Gaia18cdj, and 70 days for AT2021lwx, where these are the weighted average of the lags between the and bands and the optical band. The lags are generally longer than the lags as we would expect since the emission will include contributions from cooler, more distant dust. The width of the smoothing top hats were roughly twice the mean lag, which is the lag distribution of a spherical shell. This suggests that the dust subtends a large solid angle around the transient.

### v2

@@ Acknowledgements
https://www.cosmos.esa.int/gaia. The Gaia archive website is https://archives.esac.esa.int/gaia. This work is also based in part on observations obtained with the Samuel Oschin Telescope 48-inch and the 60-inch Telescope at the Palomar Observatory as part of the Zwicky Transient Facility project. ZTF is supported by the National Science Foundation under Grant No. AST-2034437 and a collaboration including Caltech, IPAC, the Weizmann Institute for Science, the Oskar Klein Center at Stockholm University, the University of Maryland, Deutsches Elektronen-Synchrotron and Humboldt University, the TANGO Consortium of Taiwan, the University of Wisconsin at Milwaukee, Trinity College Dublin, Lawrence Livermore National Laboratories, and IN2P3, France. Operations are conducted by COO, IPAC, and UW. The ZTF forced-photometry service was funded under the Heising-Simons Foundation grant #12540303 (PI: Graham).
@@
Funding: JTH and BJS are supported by NASA grant 80NSSC23K1431. KA is supported by the Australian Research Council Discovery Early Career Researcher Award (DECRA) through project number DE230101069. CSK is supported by NSF grants AST-1907570 and AST-2307385. CA acknowledges support by NASA grants JWSTGO-02114, JWST-GO-02122, JWST-GO-03726, JWSTGO-04436, and JWST-GO-04522. AD and BJS are supported by National Science Foundation grant AST-1911074. WBH is supported by the National Science Foundation Graduate Research Fellowship Program under Grant No. 2236415. Any opinions, findings, and conclusions or recommendations expressed in this material are those of the author(s) and do not necessarily reflect the views of the National Science Foundation.
Author contributions: JTH led the analysis and drafting the manuscript. BJS, CSK, KA, and CA helped prepare the manuscript. KA reduced and analyzed the X-ray data. CSK and JMMN conducted modeling of the IR lags. AP, JS, and TW-SH obtained follow-up spectroscopy of the sources. JTH, BJS, CSK, and JS contributed to the physical interpretation. JTH, MEH, MAT, CA, TdJ, DDD, AD, WBH, and AVP contributed to the SCAT vetting of sources. All authors provided feedback on the manuscript.
Competing interests: The authors declare no competing interests.
Data Availability: All survey photometry is publicly available. The follow-up spectra from this work are available at the Weizmann Interactive Supernova Data Repository (WISeREP, https://www.wiserep.org/). All other data needed to evaluate the conclusions in the paper are present in the paper and/or the Supplementary Materials.
List of Supplementary Materials:
Supplementary Text
Figures S1 through S8
Tables S1 through S3
References 179 through 201
@@
Figures S1 through S8
Tables S1 through S3
References 179 through 201
@@
Sample Selection
@@
The three transients we study in detail are Gaia16aaw, Gaia18cdj, and ZTF20abrbeie. Gaia16aaw, (04:11:57.000, 42:05:30.80), was discovered on 2016 January 23.5 by the Gaia Alerts team (?). The discovery was announced publicly on the Transient Name Server (TNS) and given the identification AT2016dbs (?). Gaia18cdj, (02:09:48.140, 42:04:37.02), was discovered on 2018 August 12.1 by the Gaia Alerts team (?) and announced on TNS with the identification AT2018fbb (?). ZTF20abrbeie, (21:13:48.408, 27:25:50.48), was discovered on 2021 April 13.5 by ZTF. It was announced to TNS with the identification AT2021lwx (?). Color images created using aplpy (?) from images taken by either the Dark Energy Survey (DES; for Gaia16aaw and Gaia18cdj) (?) or the Panoramic Survey Telescope and Rapid Response System (Pan-STARRS; for AT2021lwx) (?) are shown in figure S1.
As ZTF20abrbeie has had previous papers published on its evolution under the name AT2021lwx (?, ?), we elect to use that name in this manuscript to avoid confusion. For the two Gaia sources, we elect to use their survey names in this manuscript.
Archival and Transient Photometry
We first searched for available archival photometry to constrain the evolution of these objects prior to the ENT flare. For Gaia16aaw and Gaia18cdj, we obtained -band photometry from the CRTS. As the typical signal-to-noise ratio (S/N) per CRTS epoch was low, we binned these data in monthly bins to search for pre-flare variability. These data are shown in Figure 1 as tan squares. The sources are weakly detected and show no significant evidence for strong variability prior to the flares as any pre-flare variability in CRTS data for Gaia16aaw and Gaia18cdj is consistent with noise.
@@
For AT2021lwx, there is no persistent source detected in the NEOWISE single exposure catalog, consistent with the apparent lack of a host galaxy in Pan-STARRS imaging (?, ?). We therefore performed aperture photometry on the NEOWISE images to obtain our WISE light curves. To avoid contamination by nearby, bright stars, we used a 4” radius aperture with an aperture correction estimated from the computed WISE curve of growth. The local background was estimated using a sigma-clipped median within an annulus around the target. Despite the small source aperture, the band retained a low level of residual flux from the nearby, bright star. To mitigate this, we subtracted the mean pre-flare flux from the light curves and added the pre-flare scatter in quadrature with the estimated flux uncertainties for each epoch. The WISE light curves are shown in Figure 1 as olive and brown diamonds for and respectively. Given the lack of a pre-flare detection, we only show NEOWISE data for AT2021lwx.
@@
Finally, we acquired light curves from the discovering survey for each source. For Gaia16aaw and Gaia18cdj, this was -band photometry from Gaia, obtained through the Gaia Alerts service. We estimated the uncertainties on this photometry following (?) and binned visits within 5 days for subsequent analyses. The Gaia light curves are shown in Figure 1 as green hexagons. For AT2021lwx, we procured ZTF photometry in the and bands from the ZTF forced photometry service (?). We stacked the pre-flare data in monthly bins to obtain deep limits. During the rise and through peak, we stacked the data 5-day bins, switching to monthly bins after the last seasonal break to better sample the decline. These data are shown in Figure 1 as teal and red octagons.
@@
Stellar Mass and Star Formation Rate
@@
We further examine the SFRs of these ENT host galaxies using narrow emission lines and standard scaling relations. Gaia16aaw shows an emission feature near [O II], although there is contamination from night sky lines. Assuming that this line is indeed [O II], we derive a flux of erg s-1 cm-2. At the distance of Gaia16aaw and correcting for the mag extinction from the CIGALE fits, this corresponds to a luminosity of erg s-1. From typical scaling relations (?, ?), such a luminosity implies a SFR of M⊙ yr-1, roughly 2 below the estimate from the CIGALE fits.
@@
Gaia18cdj does not show an [O II] emission line, so we calculated an upper-limit on the [O II] luminosity using Equation (1) with a line width of 300 km s-1. Correcting for the extinction from the CIGALE fits, this yields a luminosity upper-limit of erg s-1, implying an SFR of M⊙ yr-1 (?, ?). While the SFRs estimated from spectroscopic scaling relations are low relative to the estimates from CIGALE, we note (1) that these lines are in regions of the spectra with S/N below 5 and (2) that the SFRs estimated from the rest-frame UV absolute magnitudes of mag (?) are consistent with the CIGALE estimates.
Dust Covering Fraction
@@
Anisotropies in the dust distribution can cause the ratio of IR to UV/optical luminosity to underestimate low covering factors and overestimate high covering factors. While the emitting geometry of these ENTs is unknown, we apply the corrections of (?) to examine their effects. Using Table 1 of (?) and assuming an aligned disk and torus with , their minimum computed optical depth, we find covering fractions of , , and for Gaia16aaw, Gaia18cdj, and AT2021lwx respectively. While these corrections increase the covering fraction estimates for our ENTs, their position relative to the TDE and ANT samples as well as the typical AGN trend in Fig. 8 is not dramatically different.
@@
We also used JAVELIN (?, ?) to compute temporal lags between the optical and IR flares, employing a top hat smoothing function. For AT2021lwx, as the source is undetected pre-flare, we added artificial optical measurements at the time of the NEOWISE observations with zero flux and uncertainties equal to the first optical flux uncertainty. This was necessary to avoid significantly negative fluxes in the model optical light curve at early times. These fits are shown in figure S5. We find rest-frame lags of 265 days for Gaia16aaw, 150 days for Gaia18cdj, and 70 days for AT2021lwx, where these are the weighted average of the lags between the and bands and the optical band. The lags are generally longer than the lags as we would expect since the emission will include contributions from cooler, more distant dust. The width of the smoothing top hats were roughly twice the mean lag, which is the lag distribution of a spherical shell. This suggests that the dust subtends a large solid angle around the transient.